The following reaction has an activation energy of
Answer #1

ln(k1/k2) = Ea/R (1/T2 - 1/T1)

Ea = 262 kJ/mol = 262000 J/mol

k1 = 6.1 x 10-8 s-1 T1 = 600 K

k2 = ? T2 = 870 K

R = 8.314 J/mol K

ln(6.1 x 10-8) - lnk2 = (262000 J/mol) / (8.314 J/mol K) [(1/870)-(1/600)]/K

-16.6124 - lnk2 = 31513 x -0.00051724

lnk2 = 16.29978-16.6124

lnk2 = -0.3126

k2 = e-0.3126

k2 = 0.73